# Deconfined quantum critical points in fermionic systems with spin-charge separation

### v3

@@ Appendix A Field theory approach
[…]
@@
The bosonization procedure amounts to rewriting the above fields as appropriate exponentials of bosonic fields $\phi_{\chi\sigma}(x)$, multiplied by a Klein factor $\eta_{\chi\sigma}$ to reproduce their correct anticommutation algebra: $\psi_{\chi\sigma}(x)=\frac{\eta_{\chi\sigma}}{\sqrt{2\pi\alpha}}e^{i\sqrt{4\pi}\chi\phi_{\chi\sigma}(x)}$. It is then customary to introduce first the linear combinations $\phi_{\sigma}(x)=\phi_{R\sigma}(x)+\phi_{L\sigma}(x)$ and $\theta_{\sigma}(x)=\phi_{R\sigma}(x)-\phi_{L\sigma}(x)$; and finally their superpositions $\varphi_{c}(x)=\phi_{\uparrow}+\phi_{\downarrow}$, and $\varphi_{s}(x)=\phi_{\uparrow}-\phi_{\downarrow}$, and similarly $\vartheta_{c}=\theta_{\uparrow}+\theta_{\downarrow}$ and $\vartheta_{s}=\theta_{\uparrow}-\theta_{\downarrow}$ At half filling $k_{F}=\pi/(2a)$ the resulting mapping reads:
@@
where, with an abuse of notation, $\chi$ and $\sigma$ in the exponent now stand for signs: $R=+$ and $L=-$, =$\uparrow$=+↑ = + and $\downarrow$=-↓ = -. Moreover $\alpha\sim a$ is an ultraviolet cutoff, and appropriate commutation relations hold between the bosonic fields. The procedure outlined above allows describing the low energy behavior of most one dimensional fermionic systems. In case of conserved total magnetization and particle number a fermionic system is described by two decoupled sine-Gordon models ${\cal H}=\sum_{\nu=c,s}{\cal H}_{\nu}^{SG}$ in the spin ($s$) and charge ($c$) degrees of freedom. Explicitly:
@@
Notice that here $v_{\nu}$, $K_{\nu}$ and $g_{\nu}$ are the excitation velocities, Luttinger parameters, and coupling amplitudes respectively, which depend on the microscopic Hamiltonian parameters.
@@ A.1 Derivation of gapped phases and symmetry breaking
In Eq. (23), the competition between the quadratic and the cosine terms determines the properties of the system. Specifically, in the first line fluctuations of the fields are promoted, whereas the cosine term in the second line favors the pinning of the field $\varphi_{\nu}(x)$ to a $x$ independent value minimizing the system energy. The resulting phase diagram can be established by means of a well known RG analysis  Alexander O. Gogolin (1998) of the model in Eq. (23). In each channel the cosine term turns out to be irrelevant if the coupling $g_{\nu}$ satisfies
@@
When the cosine term becomes relevant instead, $\varphi_{\nu}(x)$ prefers to pin to an appropriate constant value, and a gap $\Delta_{\nu}$ opens in the corresponding $\nu$-channel. Since the gapless excitations are suppressed, the long wavelength fluctuations and the associated power law behavior of correlations captured by $K_{\nu}$ are also suppressed: $K_{\nu}$ effectively flows to 00 in the $\nu$-channel.
In general, depending on the negative or positive sign of $g_{\nu}$ in Eq. (24), the pinning values of $\varphi_{\nu}$ are respectively 00 and $\sqrt{\pi/8}$. In presence of total charge/spin vector conservation of the microscopic lattice Hamiltonian though, the RG flow in the corresponding channel is restricted to solely the line $2\pi v_{\nu}(K_{\nu}-1)=g_{\nu}$, implying that the cosine term can become relevant only for $g_{\nu}<0$ ($\varphi_{\nu}=0$). In this case $K_{\nu}=1+\frac{g_{\nu}}{2\pi v_{\nu}}$: the theory is gapless for $K_{\nu}>1$ and gapped for $K_{\nu}<1$. More generally, rooted in the total charge/spin conservation of the lattice Hamiltonian is the global $U(1)$ symmetry
@@
so that the full symmetry in a gapless channel will be enriched to $U(1)\times U(1)\cong SU(2)$  555notice that the effective $SU(2)$ symmetry holds in the case of $SU(2)$ symmetric choices of the microscopic parameters of a Hamiltonian. When the cosine term becomes relevant though, the above symmetry becomes discrete, namely
@@
Relevantly, also other discrete like the translational, spin-flip and inversion symmetries characterize ${\cal H}$, see for instance  Nakamura (2000); Mudry et al. (2019). According to the Landau’s picture  Landau et al. (1999), some of these could be spontaneously broken generating LO phases captured by a corresponding local order parameter.
@@ A.2 Bosonization of the microscopic model
$H=$displaystyle-- $t\sum_{j,\sigma}(c_{j,\sigma}^{\dagger}c_{j+1,\sigma}+h.c.)+U\sum_{j}n_{j,\uparrow}n_{j,\downarrow}$ (28)
@@ Appendix B Extrapolation of order parameters around criticality
To address the behavior of the order parameters around the DQCPs in Sections III.2.1 and III.2.2, we have to properly extrapolate their behavior in the limit of infinite bond dimension $\chi\rightarrow\infty$. This is crucial as the scaling of these quantities around the critical point leads us to obtain critical exponents incompatible with a transition in the Ising universality class. This, in turn, corroborates our conclusion that we indeed find a DQCP. However, the convergence criterion usually used in VUMPS, only assures that we find a variational minimum in the manifold of the finite bond dimension ansatz in consideration Zauner-Stauber et al. (2018). This does not readily give a measure of the quality of the approximation when compared to the physical ground state.
To overcome this problem we proceed as in standard DMRG, by performing an extrapolation as a function of the discarded Schmidt values after performing a truncation during the optimization process Schollwöck (2011).
In particular, to obtain this truncation error, it is sufficient to apply the Hamiltonian $H$ projected on the subspace of the tensors currently being optimized, obtaining a state of bond dimension $d\chi$, where $d=4$ in the case of spinful fermions. We can then obtain the new ground state in this subspace by performing an exact diagonalization. Finally, by truncating this state back to bond dimension $\chi$ we can obtain the error. This procedure is straightforward in the case of two-site DMRG, where the growing of the bond dimension is dynamically performed during the optimization sweep. In the case of VUMPS, however, the bond dimension is kept fixed during the optimization process. We then proceed in the following way: (i) we perform a subspace expansion to express the converged ground state at bond dimension $\chi$ using a larger bond dimension $4\chi$ Hubig et al. (2015). (ii) We perform one sweep of the VUMPS optimization, equivalent to a single application of $H$ to the unit cell and an exact diagonalization in this subspace, (iii) we manually compute the error
@@
where $S_{i}$ are the Schmidt values along the bond on which the Hamiltonian is applied. This procedure follows the one explained in Appendix A4 of Zauner-Stauber et al. (2018).

### v4

@@ Appendix A Field theory approach
The standard field theory treatment of 1D lattice systems of correlated fermions is based on a weak coupling approach, known as bosonization. As this technique has been largely employed, see  [48, 49] for exhaustive references, here we just summarize the relevant steps needed to derive the low energy properties of a many-body fermionic system with short range interaction within Abelian bosonization. The starting point consists in defining the continuum limit by replacing the discrete sum over sites $j$ with integrals over the coordinate $x=ja$, i.e. $\sum_{j}\rightarrow\frac{1}{a}\int dx$ with $a$ lattice spacing. Thanks to the peculiar nature of the one dimensional Fermi surface consisting of just two disconnected points $\pm k_{F}$, usual fermionic operators $c_{j\sigma},c_{j\sigma}^{\dagger}$, with $\sigma=\uparrow,\downarrow$ labeling the spin orientation, can be rewritten in terms of right ($R$) and left ($L$) fermionic fields $\psi_{\chi\sigma}(x)$ ($\chi=R,L$),
@@
The bosonization procedure amounts to rewriting the above fields as appropriate exponentials of bosonic fields $\phi_{\chi\sigma}(x)$, multiplied by a Klein factor $\eta_{\chi\sigma}$ to reproduce their correct anticommutation algebra: $\psi_{\chi\sigma}(x)=\frac{\eta_{\chi\sigma}}{\sqrt{2\pi\alpha}}e^{i\sqrt{4\pi}\chi\phi_{\chi\sigma}(x)}$. It is then customary to introduce first the linear combinations $\phi_{\sigma}(x)=\phi_{R\sigma}(x)+\phi_{L\sigma}(x)$ and $\theta_{\sigma}(x)=\phi_{R\sigma}(x)-\phi_{L\sigma}(x)$; and finally their superpositions $\varphi_{c}(x)=\phi_{\uparrow}+\phi_{\downarrow}$, and $\varphi_{s}(x)=\phi_{\uparrow}-\phi_{\downarrow}$, and similarly $\vartheta_{c}=\theta_{\uparrow}+\theta_{\downarrow}$ and $\vartheta_{s}=\theta_{\uparrow}-\theta_{\downarrow}$ . At half filling $k_{F}=\pi/(2a)$ the resulting mapping reads:
@@
where, with an abuse of notation, $\chi$ and $\sigma$ in the exponent now stand for signs: $R=+$ and $L=$-, $\uparrow=+$ and $\downarrow=$-. Moreover $\alpha\sim a$ is an ultraviolet cutoff, and appropriate commutation relations hold between the bosonic fields. The procedure outlined above allows describing the low energy behavior of most one dimensional fermionic systems. In case of conserved total magnetization and particle number a fermionic system is described by two decoupled sine-Gordon models ${\cal H}=\sum_{\nu=c,s}{\cal H}_{\nu}^{SG}$ in the spin ($s$) and charge ($c$) degrees of freedom. Explicitly:
@@
Notice that here $v_{\nu}$, $K_{\nu}$ and $g_{\nu}$ are the excitation velocities, Luttinger parameters, and coupling amplitudes respectively, which depend on the microscopic Hamiltonian parameters. Other cosine terms, which may appear within each sine-Gordon Hamiltonian ${\cal H}_{\nu}^{SG}$ or as coupling terms between the charge and spin sectors, are typically neglected in the low-energy theory. This is because they involve faster oscillations in the bosonic fields and, accordingly, acquire larger scaling dimensions. From a renormalization group (RG) perspective, such terms are therefore irrelevant and do not affect the universal low-energy physics.
@@ A.1 Derivation of gapped phases and symmetry breaking
In Eq. (24), the competition between the quadratic and the cosine terms determines the properties of the system. Specifically, in the first line fluctuations of the fields are promoted, whereas the cosine term in the second line favors the pinning of the field $\varphi_{\nu}(x)$ to a $x$ independent value minimizing the system energy. The resulting phase diagram can be established by means of a well known RG analysis  [48] of the model in Eq. (24). In each channel the cosine term turns out to be irrelevant if the coupling $g_{\nu}$ satisfies
@@
When the cosine term becomes relevant instead, $\varphi_{\nu}(x)$ prefers to pin to an appropriate constant value, and a gap $\Delta_{\nu}$ opens in the corresponding $\nu$-channel. Since the gapless excitations are suppressed, the long wavelength fluctuations and the associated power law behavior of correlations captured by $K_{\nu}$ are also suppressed: $K_{\nu}$ effectively flows to 0 in the $\nu$-channel.
In general, depending on the negative or positive sign of $g_{\nu}$ in Eq. (25), the pinning values of $\varphi_{\nu}$ are respectively 0 and $\sqrt{\pi/8}$. In presence of total charge/spin vector conservation of the microscopic lattice Hamiltonian though, the RG flow in the corresponding channel is restricted to solely the line $2\pi v_{\nu}(K_{\nu}-1)=g_{\nu}$, implying that the cosine term can become relevant only for $g_{\nu}<0$ ($\varphi_{\nu}=0$). In this case $K_{\nu}=1+\frac{g_{\nu}}{2\pi v_{\nu}}$: the theory is gapless for $K_{\nu}>1$ and gapped for $K_{\nu}<1$. More generally, rooted in the total charge/spin conservation of the lattice Hamiltonian is the global $U(1)$ symmetry
@@
so that the full symmetry in a gapless channel will be enriched to $U(1)\times U(1)\cong SU(2)$  666notice that the effective $SU(2)$ symmetry holds in the case of $SU(2)$ symmetric choices of the microscopic parameters of a Hamiltonian. When the cosine term becomes relevant though, the above symmetry becomes discrete, namely
@@
[…]
@@ A.2 Bosonization of the microscopic model
$H=$ - $t\sum_{j,\sigma}(c_{j,\sigma}^{\dagger}c_{j+1,\sigma}+h.c.)+U\sum_{j}n_{j,\uparrow}n_{j,\downarrow}$ (29)
@@ Appendix B Extrapolation of order parameters around criticality
To address the behavior of the order parameters around the DQCPs in Sections III.2.1 and III.2.2, we have to properly extrapolate their behavior in the limit of infinite bond dimension $\chi\rightarrow\infty$. This is crucial as the scaling of these quantities around the critical point leads us to obtain critical exponents incompatible with a transition in the Ising universality class. This, in turn, corroborates our conclusion that we indeed find a DQCP. However, the convergence criterion usually used in VUMPS, only assures that we find a variational minimum in the manifold of the finite bond dimension ansatz in consideration [51]. This does not readily give a measure of the quality of the approximation when compared to the physical ground state.
To overcome this problem we proceed as in standard DMRG, by performing an extrapolation as a function of the discarded Schmidt values after performing a truncation during the optimization process [113].
In particular, to obtain this truncation error, it is sufficient to apply the Hamiltonian $H$ projected on the subspace of the tensors currently being optimized, obtaining a state of bond dimension $d\chi$, where $d=4$ in the case of spinful fermions. We can then obtain the new ground state in this subspace by performing an exact diagonalization. Finally, by truncating this state back to bond dimension $\chi$ we can obtain the error. This procedure is straightforward in the case of two-site DMRG, where the growing of the bond dimension is dynamically performed during the optimization sweep. In the case of VUMPS, however, the bond dimension is kept fixed during the optimization process. We then proceed in the following way: (i) we perform a subspace expansion to express the converged ground state at bond dimension $\chi$ using a larger bond dimension $4\chi$ [114]. (ii) We perform one sweep of the VUMPS optimization, equivalent to a single application of $H$ to the unit cell and an exact diagonalization in this subspace, (iii) we manually compute the error
@@
where $S_{i}$ are the Schmidt values along the bond on which the Hamiltonian is applied. This procedure follows the one explained in Appendix A4 of [51].